OpenStudy (anonymous):

I'm having a hard time sketching the trig function graphs. Can someone make it more clear? My first problem is ** Sketch the graph of y=tan(pix/4), where x is measured in radians. In your graph, show at least one full period, and label asymptotes.

OpenStudy (welshfella):

first thing:- when x = 0 then tan (pix/4) is also 0 so the graph passes through the origin Also when x = pi/2 (90 degrees) there is no value for tan so there is a vertical asymptote at x = pi/2. This will also occur at x = -pi/2 , 3pi/2 and -3pi/2

OpenStudy (welshfella):

Mark the x axis in units of pi/4

OpenStudy (anonymous):

Is it always units pi/ 4?